OpenStudy (anonymous):

Solve the proportion: 16/17=x/40

OpenStudy (anonymous):

first step: cross multiply second step divide by 17 third step : calculate!

OpenStudy (anonymous):

\[\LARGE \frac{16}{17}=\frac{x}{40} \quad \quad \Longrightarrow 17x=40\cdot 16\] this was the first step... do the rest :)

OpenStudy (anonymous):

this is not really helping you knoee..

OpenStudy (anonymous):

17x=640 then divide 17 by both side right? then you get x= ?

OpenStudy (anonymous):

\[\LARGE \frac{17x}{17}=\frac{640}{17} \] \[\LARGE \frac{17}{17}\cdot x=\frac{640}{17} \] \[\LARGE x=\frac{640}{17} \] or \[\LARGE x\approx 37.64\]

OpenStudy (anonymous):

i was correct the answer is 37.6

OpenStudy (anonymous):

I don't see where you wrote that, and I denied !! O_O ...